# Uncertainty-Guided Optimization on Large Language Model Search Trees

### v1

@@ 5.2.3 Runtime
Finally, we analyze the runtime overhead on top of the LLM’s forward pass due to ULTS. Figure 6 shows the wall-clock time per iteration, averaged over all sentences and all iterations, broken down into the time spent on the LLM’s forward pass and the time spent on sampling the optimal values and optimizing the acquisition function. The error bars indicate the $95\%$- confidence intervals. Results are shown for the experiments with $k_{\text{max}}=20$ and the Dirichlet prior—the empirical prior performs similarly since ULTS does not differentiate between them in Alg. 2. We note that the runtime overhead of ULTS is small compared to the time spent to do an LLM forward pass. However, note that our current implementation only expands one node of the search tree in each iteration and thereby only evaluate one token sequence per forward-pass. Depending on the available hardware, it may be advantageous to do batch node expansions. [^3] The memory requirement for a forward pass depends on the model size, sequence length and batch size.
@@ 6 Conclusion
We have discussed our method, Uncertainty-guided Likelihood-Tree Search (ULTS), a probabilistic algorithm for the decoding process of large-language-model (LLM) search trees. Our method quantifies and leverages computational uncertainty over the maximum value of the optimization process. ULTS exploits the structure of the optimization problem—in particular, the concentration strength of the LLM softmax outputs—by putting a prior over the LLM’s softmax outputs, and use the implied posterior samples to guide exploration-exploitation in a non-myopic manner. It improves on the search efficiency of standard beam search by adapting the number of nodes expanded per level of the tree. In particular, ULTS expands fewer nodes than beam search, but achieves higher likelihood on average. As an interesting future direction, one can study the effect of using more sophisticated priors, e.g. where iid. is not assumed. One can also consider batched acquisition strategies, similar to batch Bayesian optimization techniques [7, 43]. Moreover, it is also interesting to take into account the uncertainty over the LLM’s outputs (e.g. in the context of Bayesian LLMs), and to incorporate (possibly uncertain) external rewards (such as ones coming from an RLHF-trained reward model) in addition to the current likelihood rewards.
@@ A.1 Prior assumptions
The symmetry assumption for the Dirichlet prior on the LLM’s softmax outputs likely does not entirely hold—some words occur more often in natural language than others [16]. However, ULTS only requires access to the distribution over the maximum of the unexplored part of the search space and not over the arg max. The former is not affected by permutations of the entries in the categorical distributions, which is why we suspect that a symmetric Dirichlet distribution with sufficiently small concentration parameter is a good proxy. Moreover, the empirical prior can also be used to address this limitation, without incurring large overhead.
@@ Appendix B Additional Experimental Details
For the additional machine translation experiment, we use the WMT-19 German to English dataset https://huggingface.co/facebook/wmt19-de-en and the following Llama-2-7b based translation model https://huggingface.co/kaitchup/Llama-2-7b-mt-German-to-English.
@@ C.1 On-model experiments
Figure 7 shows the additional results from the experiment on on-model samples from Dirichlet priors for $\alpha=0.5$ and $\alpha=0.8$ as described in Section 4.1 of the main text. Here, too, the probabilistic versions outperform beam search.
Figure 7: Comparison on trees, where the transition probabilities are sampled from a Dirichlet prior for two different values of the concentration parameter.
@@ C.2 Machine translation
We compare ULTS and beam search in a machine translation task with 1000 randomly sampled sequences from the standard WMT-19 German to English dataset in Fig. 8. Unlike other settings we have used in the main text, here we terminate exploration once an <EOS> token is found. This setting is more ubiquitous in practice, but harder to compare in terms of likelihood due to the different sentence lengths. We observe that ULTS with the Dirichlet prior with $\alpha=0.0001$ underperform compared to beam search. This is because this prior induces less exploration. Indeed, when $\alpha$ is reduced, ULTS can utilize the search budget much more efficiently, achieving significantly better likelihoods while invoking much fewer LLM calls than beam search.
Figure 8: Machine translation results with the WMT-19 German-to-English dataset. All algorithms stop at the <EOS> token.
C.3 Alternative acquisition function

### v2

@@ 5.3 Runtime
We analyze the runtime overhead on top of the LLM’s forward pass due to ULTS. Figure 5 shows the wall-clock time per iteration, averaged over all sentences and all iterations, broken down into the time spent on the LLM’s forward pass and the time spent on sampling the optimal values and optimizing the acquisition function. The error bars indicate the $95\%$- confidence intervals. Results are shown for the experiments with $k_{\text{max}}=20$ and the Dirichlet prior—the empirical prior performs similarly since ULTS does not differentiate between them in Alg. 2. We note that the runtime overhead of ULTS is small compared to the time spent to do an LLM forward pass. However, note that our current implementation only expands one node of the search tree in each iteration and thereby only evaluate one token sequence per forward-pass. ULTS can be extended similar like BO can be extended into batch BO. It outside of the present work’s scope but is a promising future work.
@@
5.4 Extensions
Method Perplexity $\downarrow$ Diversity $\uparrow$
Beam search 1.16$\pm$0.01 0.33$\pm$0.01
Nucleus sampling 3.69$\pm$0.07 0.82$\pm$0.01
Contrastive search 1.27$\pm$0.01 0.4$\pm$0.01
ULTS 1.16$\pm$0.01 0.33$\pm$0.01
ULTS with $\tilde{u}$ 1.3$\pm$0.01 0.55$\pm$0.01
Table 1: Different utility functions (here, $\tilde{u}$) can be used in ULTS’ acquisition function. E.g., to enforce low perplexity and high diversity.
Let ${\bm{v}}_{i}$ contains the optimal values of the descendants of $x_{i}$ and let $u(x_{c},{\bm{v}}_{i}):=\mathbb{I}[v_{\hat{x}_{c}}=\max_{x_{j}\in\mathrm{children}(x_{i})}v_{\hat{x}_{j}}]$. The acquisition function (1) we employed so far can be seen as a sample-based approximation to the expectation
[…]
From the Bayesian decision-theoretic perspective, $u(x_{c},{\bm{v}}_{i})$ can thus be seen as a utility function, encoding our preferences when making decisions under the posterior belief. Hence, we can replace it with any other utility function just like in standard Bayesian optimization (Garnett, 2023) and then use the posterior samples to derive the acquisition function (Wilson et al., 2018).
For example, we can use a utility function that encodes our preference of picking sequences with minimal repetition, which is desirable to avoid text degeneration (Holtzman et al., 2019). Concretely, we can employ a new utility function that contains a repetition-penalty term: $\tilde{u}(x_{c},{\bm{v}}_{i})=\mathbb{I}[(\log v_{\hat{x}_{c}}+\lambda\,b(\hat{x}_{c}))=\max_{x_{j}\in\mathrm{children}(x_{i})}(\log v_{\hat{x}_{j}}+\lambda\,b(\hat{x}_{j}))]$, where $b(\hat{x}_{c})$ is a log-diversity term (Su et al., 2022). Details in Section C.7. The results of this new utility function is in Table 1: ULTS can achieve good perplexity and diversity at the same time, showing the flexibility of our probabilistic decision-theoretic perspective. More results can be found in Table 3 in Section C.7.
@@ 6 Conclusion
We have discussed our method, ULTS, a probabilistic decision-making algorithm for the decoding process of large-language-model (LLM) search trees. Our method quantifies and leverages computational uncertainty over the maximum value of the optimization process. ULTS exploits the structure of the optimization problem—in particular, the concentration strength of the LLM softmax outputs—by putting a prior over the LLM’s softmax outputs and using the implied posterior samples to guide exploration-exploitation in a non-myopic manner. Our work thus connects two seemingly disjoint fields: Bayesian optimization and natural language processing, opening up opportunities for interesting future work in probabilistic inference for LLMs.
Future work  One can study the effect of using more sophisticated priors, e.g. where iid. is not assumed. One can also consider batched acquisition strategies, similar to batch Bayesian optimization techniques (González et al., 2016; Wu & Frazier, 2016). Moreover, it is also interesting to incorporate the uncertainty over the LLM’s outputs (e.g. in the context of Bayesian LLMs) in order to account for possible miscalibration of the likelihood. Another research direction is to extend the approach to reasoning tasks beyond language tasks, e.g. by including beliefs over external rewards (Zhang et al., 2023), such as ones coming from an RLHF-trained reward model, in addition to the current likelihood rewards. Finally, our work in connecting Bayesian optimization with LLMs paves the way to perform probabilistic reasoning with the tree of thoughts framework (Yao et al., 2023).
@@ A.1 Prior assumptions
The symmetry assumption for the Dirichlet prior on the LLM’s softmax outputs likely does not entirely hold—some words occur more often in natural language than others (Kingsley Zipf, 1932). However, ULTS only requires access to the distribution over the maximum of the unexplored part of the search space and not over the arg max. The former is not affected by permutations of the entries in the categorical distributions, which is why we suspect that a symmetric Dirichlet distribution with sufficiently small concentration parameter is a good proxy. Moreover, the empirical prior can also be used to address this limitation, without incurring large overhead.
@@ C.1 On-model experiments
We compare ULTS to beam search on artificially generated search problems from Dirichlet priors. The trees have branching factor $b=8$ and depth $d=5$. Since these trees are so small we optimized the acquisition function in eq. (1) over the the full boundary, i.e. non-recursively. The transition probabilities at each node are sampled from a Dirichlet prior with fixed $\alpha\in\{0.1,0.2,0.5,0.8\}$. The comparison is on-model, i.e., ULTS is run with the ground truth parameter of $\alpha$. We repeat the experiment with different values for the confidence parameter $\varepsilon$ of ULTS from $\{0.05,0.1,0.3\}$. Since the toy problems are so small, the exploration of too many nodes is not an issue and we use $k_{\max}=\infty$. Beam search is run with beam sizes ranging from $1$ to $7$. The results in Fig. 6 show that ULTS dominates across the entire range of hyper-parameters. This suggests that knowledge about concentration strength helps reduce the number of search steps.
Figure 6: Comparison on trees, where the transition probabilities are sampled from a Dirichlet prior for different values of the concentration parameter.
@@ C.2 Machine translation
We show some example sequences found with ULTS and beam search. For these examples, we picked sentences where ULTS achieves better BLEU scores. Other examples have mainly the same outputs as beam search while ULTS is more efficient.
Example 1 input prompt: It is annoying when geographical maps are not up-to-date.
ground truth: Es nervt, wenn Landkarten nicht aktuell sind.
ULTS translation: Es ist ärgerlich, wenn geographische Karten nicht aktuell sind.
ULTS BLEU/log-likelihood: 0.189/-3.914
Beam search translation: Es ist ärgerlich, wenn die geographischen Karten nicht auf dem neuesten Stand sind.
Beam search BLEU/log-likelihood: 0.0/-4.230
Example 2 input prompt: The historical maps of the digital BayernAtlas, an offering from the State Government’s Geoportal Bayern, are anything but up-to-date – and yet it is precisely for this reason that they are so informative.
ground truth: Die historischen Landkarten des digitalen Bayern-Atlases, ein Angebot des Geoportals Bayern der Staatsregierung, sind alles andere als aktuell - doch gerade deshalb sehr aufschlussreich.
ULTS translation: Die historischen Karten des digitalen BayernAtlas, ein Angebot des Landesgeoportals Bayern, sind alles andere als aktuell – und gerade deshalb so informativ.
ULTS BLEU/log-likelihood: 0.292/-9.247
Beam search translation: Die historischen Karten des digitalen BayernAtlas, ein Angebot des Landesgeoportals Bayern, sind alles andere als aktuell – und gerade deshalb sind sie so informativ.
Beam search BLEU/log-likelihood: 0.272/-9.59
Example 3 input prompt: Even if the French troops finally retreated with the Treaty of Lunéville from 9th February 1801: it was the current neighbours who had the idea to create a comprehensive map of Bavaria.
ground truth: Auch wenn die französischen Truppen mit dem Frieden von Lunéville vom 9. Februar 1801 schließlich abzogen: Es waren die heutigen Nachbarn, die die Idee einer flächendeckenden Bayern-Karte kreierten.
ULTS translation: Auch wenn sich die französischen Truppen mit dem Vertrag von Lunéville vom 9. Februar 1801 schließlich zurückzogen: Es waren die heutigen Nachbarn, die die Idee hatten, eine umfassende Karte von Bayern zu erstellen.
ULTS BLEU/log-likelihood: 0.458/-13.111
Beam search translation: Auch wenn sich die französischen Truppen schließlich mit dem Vertrag von Lunéville vom 9. Februar 1801 zurückzogen: Es waren die heutigen Nachbarn, die die Idee hatten, eine umfassende Karte von Bayern zu erstellen.
Beam search BLEU/log-likelihood: 0.400/-12.754
C.3 Different choice of hyperparameter $\alpha$ for the Dirichlet prior
For the Dirichlet prior, indeed $\alpha$ has a meaningful impact on the performance of ULTS: It can be seen as a hyperparameter that trades exploitation for exploitation (smaller value means more exploration). So, the choice of $\alpha$ should be based on how much budget one has since more exploration means more node expansions. Figure 8 shows a comparison of two different values $\alpha=0.0001$ and $\alpha=5e-6$.
Figure 8: Comparison of two different values $\alpha=0.0001$ and $\alpha=5e-6$ of the Dirichlet prior concentration parameter in ULTS
C.4 Additional runtime results
Method $\bm{k}$ or $\bm{k_{\mathrm{max}}}$ Time in s
Beam search 5 $2.158\pm 0.029$
ULTS Dirichlet 3 $4.622\pm 0.114$
ULTS Dirichlet 5 $5.177\pm 0.164$
Table 2: Total runtime for decoding one of the TL;DR input prompts with Llama-2-7b.
In Table 2 we show runtime results for beam search with beam size $k=5$ and ULTS with maximal beam size $k_{max}=3$ and $k_{max}=5$. Despite expanding fewer nodes than beam search, ULTS is currently slowerin settings where different nodes expansion in beam search can be batched. However, note that batching is not always possible, e.g. in memeory constraint settings (the memory resources depend on the model size, sequence length, as well as batch size).
C.5 Alternative acquisition function
@@
C.6 Full results for decoding experiments with LLama-2-7b and GPT-2
Figure 11 shows the full results from the decoding experiments with LLama-2-7b and GPT-2 from section 5 in the main text.
Figure 11: Full results for decoding experiments with LLama-2-7b and GPT-2
C.7 Details for experiment with repetition penalty
[…]
$\textbf{diversity}((a_{1},...,a_{i}))=\prod_{n=2}^{4}(1-\textbf{rep-n}((a_{1},...,a_{i}))).$
Since we ULTS optimizes the total likelihood of a sequence and not the average likelihood of a sequence, we additionally scale the term with the tree depth, leading to $b(x)=d\cdot\log\textbf{diversity}((a_{1},...,a_{i}))$ in log-space. For ULTS we use a Dirichlet prior with concentration parameter $\alpha=5\times 10^{-6}$, $\epsilon=0.1$, $k_{max}=5$. and penalty parameter $\lambda\in\{0.0,0.2,0.4,0.6,0.8,1.0\}$. The parameters for contrastive search are top-$k$ $=5$ and penalty parameter $\alpha=0.6$. For beam search, we use $k=5$. The main paper reports results for ULTS with $\lambda=0.0$, i.w. without penalty, and $\lambda=0.2$. The full results are given in Table 3. One can see that with increasing penalty the diversity of the decoded sequences increases, i.e. extending the acquisition function with a repetition term is effective.
Method Perplexity $\downarrow$ Diversity $\uparrow$
Beam search 1.16$\pm$0.01 0.33$\pm$0.01
Nucleus sampling 3.69$\pm$0.07 0.82$\pm$0.01
Contrastive search 1.27$\pm$0.01 0.4$\pm$0.01
ULTS 1.16$\pm$0.01 0.33$\pm$0.01
ULTS with $\lambda=0.2$ 1.3$\pm$0.01 0.55$\pm$0.01
ULTS with $\lambda=0.4$ 1.45$\pm$0.01 0.74$\pm$0.01
ULTS with $\lambda=0.6$ 1.5$\pm$0.01 0.79$\pm$0.01
ULTS with $\lambda=0.8$ 1.53$\pm$0.01 0.82$\pm$0.01
ULTS with $\lambda=1.0$ 1.55$\pm$0.01 0.83$\pm$0.01
Table 3: Effects of a penalty term in ULTS’ acquisition function.